OpenStudy (anonymous):

Fan and Medal A town plans to make a triangular park. The triangle has a base of 120 feet and a height of 115 feet. What will the area of the park be?

OpenStudy (the_fizicx99):

Simply apply the formula, A = 1/2 * b * h A = 1/2 * 120 * 115 First start by multiply 120 * 115 then multiplying the result by 1/2 Or, divide the base by 2 and multiply it by the height. I find it easier to just divide the base by 2

OpenStudy (anonymous):

Thank you :3